OpenStudy (anonymous):

Solve for m. m/2 - √5= 3

OpenStudy (jhannybean):

\[\frac{m}{2}-\sqrt{5}=3\]Add \(+\sqrt{5}\) to both sides.

OpenStudy (jhannybean):

now you'll have \[\frac{m}{2} = 3+\sqrt{5}\]

OpenStudy (jhannybean):

Multiply both sides of the equation by 2 to isolate the m on one side. \[2\left(\frac{m}{2} = 3+\sqrt{5}\right)\]

OpenStudy (jhannybean):

\[m = 2(3+\sqrt{5})\]

OpenStudy (jhannybean):

Simplify.

OpenStudy (anonymous):

6+2√5

OpenStudy (anonymous):

6+2 sqrt5

OpenStudy (anonymous):

Right??

OpenStudy (jhannybean):

Yep thats what I got.

OpenStudy (anonymous):

thank you